For a test of Ho: p = 0.50 and Ha: p < 0.50, the sample proportion is 0.42 based on a sample size of 100. Report the value of the test statistic. Round your final answer to three decimal places and enter only a number. (That is, do not include any letter or the = sign)
Answer #1

solution

this is left tailed test

=0.42

P0 = 0.50

1 - P0 = 1 -0.50 =0.50

n=100

Test statistic = z

= - P0 / [P0 * (1 - P0 ) / n]

=0.42 -0.50 / [(0.50*0.50) / 100]

z= -1.600